OpenStudy (anonymous):

Determine the solutions to the quadratic equation -4x2 - 32x - 64 = 0.

OpenStudy (nurali):

-4x^2 - 32x - 64 = 0 -4 (x^2 + 8x + 16) = 0 -4 (x +4)(x+4) = 0 x+ 4 =0 x = -4

OpenStudy (anonymous):

Factor out -4: -4(x^2 + 8x + 16) = 0 We can take out the -4 by dividing both sides by -4 since we don't really need it to find the roots. So we get: x^2 + 8x + 16 = 0 Now we factor the quadratic. To do that we need to find two numbers which multiply to 16 and add up to 8. Such two numbers are 4 and 4. (x + 4)(x + 4) = (x + 4)^2 = 0 Can you find which x-value would satisfy the equation now that we have factored? @amandaschlor